OpenStudy (anonymous):

log 6 2/3 expand the expression

OpenStudy (anonymous):

log 6 2/3 is this log base 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\log_{6} \frac{ 2 }{ 3 }\] change of base to 10 \[=\frac{ \log_{10}\frac{ 2 }{ 3 } }{ \log_{10}6 }\]=?

OpenStudy (anonymous):

you can use your calculator ... :D

OpenStudy (anonymous):

lol i still dont know how to figure it out like srsly im really that stupid XD